OpenStudy (anonymous):

Essay Question! Need help!! 9.02] What are the solutions of x2 + 3x − 18 = 0? Show all work for full credit.

OpenStudy (anonymous):

this one is pretty simple, all you have to do is is factor it normally (x+6)(x-3)

OpenStudy (anonymous):

oh, okay, ,lol! :D

OpenStudy (anonymous):

then set each one to 0 x+6=0 x-3=0

OpenStudy (anonymous):

wait lol i wasnt done

OpenStudy (anonymous):

x=-6, 3

OpenStudy (anonymous):

I know, so that is it. Wow. Thx! LOL. :)

OpenStudy (anonymous):

np lol

OpenStudy (anonymous):

Just one more question, how does positive 6, change to negative 6? And negative 3 to positive 3?

OpenStudy (anonymous):

hen set each one to 0 x+6=0 x-3=0 this step here is the reason you subtracted/ added the number causing their signs to change

OpenStudy (anonymous):

Oh, ok, thx, lol, got it! :)